# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v6

@@ Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I digitize a unique daily longitudinal survey dataset on household budgets to examine the extent to which consumption is affected by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases with local retailers played a role in smoothing short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
@@ 2.1 Japanese Economy around the First World War
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934-1935 price. Sources: Created by the author using Ohkawa et. al. (1974, p. 237).
@@ 2.2 Tsukishima: A Representative Manufacturing Area
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative district in 1920. Geographical coordinate system is based on JGD2000/(B, L). Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
Figure 2: Administrative Wards in Tokyo City
@@
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery sector and the other manufacturing sectors, respectively. Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure 3: Factories in Tsukishima
@@
Notes: Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males. Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. Since the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals. The figures for Tsukishima listed in column 3 of both panels are based on the total sum of statistics of all the blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme). Sources: Columns 1 and 2 of Panel A use the Statistics Bureau of the Cabinet (1929c, pp. 18–19). Columns 1 and 2 of Panel B use the Statistics Bureau of the Cabinet (1929c, pp. 38–43); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 3 of Panel A uses the Tokyo City Office (1922c, pp. 42–47). Column 3 of Panel B uses the Tokyo City Office (1922a, pp. 26–31); Tokyo City Office (1922b, pp. 2–3; 42–46).
@@ 2.3 Risk-Coping Institutions
Working-class households had been exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can help mitigate the idiosyncratic income shocks on the workers. [^24] Gertler and Gruber (2002) shows the impacts of illness on consumption smoothing behavior in early 1990s Indonesia and argues the importance of public support for the illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar U.S.. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as the case of the U.S. from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The former health insurance was enacted in 1927 and unemployment insurance had not been established throughout the prewar period. Therefore, the workers had basically no public support for their illness, injuries, and joblessness. Instead, they might have used private financial institutions to deal with the idiosyncratic risks.
@@
Several financial institutions were indeed available for working-class households. A representative savings institution for households was the postal savings (yūbin chokin). In Tokyo City, roughly $1.4$ million people had postal savings accounts, covering approximately $67$% of the entire citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which is less than one month’s earnings of the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) is another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was roughly $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were the commerce and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings bank was an institution designed for lower-income workers than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it had been used among small business owners and merchants for running their businesses (Online Appendix A.2). [^25] Generally, the ordinary bank in prewar Japan was the institution for providing large loans to businesses. There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as this Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions, and transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
@@
There also be a few lending institutions. Credit purchases (kakegai) in retail shops were the most popular institution. Although the systematic statistics on credit purchases are unavailable, the THBS households often used credit purchases in various retail shops for rice, fish, vegetables, firewood, and charcoal. [^26] Cooperative societies were also available for worker households to buy daily commodities, albeit they were not lending institutions. There were 26 cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had roughly $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand. Most of the THBS households had not used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop was another lending institution. The accessibilities were higher than the other lending institutions: [^27] The number of pawnshops in Tokyo were $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. there were $13$ pawnshops in Tsukishima, more than twice the large number of moneylenders (Section 2.2). However, the core users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% of all users. In contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). In this light, the pawnshop is a small lending institution for low-income working-class households who might have faced credit constraints. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles were inexpensive clothes (Shibuya and Shojiro 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates.  [^28] This means that the average redemption rate was substantially high: it was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
@@
To summarize, while there were several institutions, the available institutions for the factory workers were few in the practical sense. For savings institutions, the postal savings might have been the best available devices for Tsukishima workers. Regarding lending institutions, credit purchase is the plausible device among factory workers. While pawnshops were accessible financial institutions, they might not provide sufficient money as the articles were usually inexpensive. The factory workers in Tsukishima might have combined these devices to deal with the idiosyncratic income shocks.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption smoothing and risk-coping behaviors requires the panel structures for aggregate income and expenditure (Section 4). Among the 40 THBS households, there are five cross-sectional units in the semi-monthly and monthly data. [^34] Section 3.3 provides details on the definitions of aggregate income and expenditure. I excluded these five cross-sectional units, leaving 35 THBS households. Next, a household with incomplete income information, which is necessary for my analysis, is removed. Finally, I exclude a household for which no information on payments to the credit purchases was recorded. [^35] Online Appendix B.3 summarizes the finer details of the original data structure and trimming. Consequently, my THBS sample includes 33 households with panel structures from 12th January to 11th November 1919.
Before assessing the representatives of the THBS sample, I test the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null results. This result is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference is less likely to be correlated with the unbalancedness.
@@
Average monthly earnings 52.7 55.0
95%CI [48.0, 57.4]
@@
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. The definition of classification in the Statistics of the National Police Agency (column 1) is used in columns 2 and 3. Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be regarded as one of the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389). Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The third to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5). Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30-40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation. Sources: Panel A: Column 1 uses the National Police Agency (1920b, p. 236). Column 2 uses the Department of Health, Ministry of the Interior (1923a, p. 385–386). Column 3 uses the Department of Health, Ministry of the Interior (1923a, p. 154). Panel B: Column 1 uses the Tokyo City Office (1922a, pp. 26–31; 262–283); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 2 uses the National Police Agency (1920b, p. 53). The figures in column 3 are calculated by the author using the THBS dataset. Panel C: Column 1 uses the THBS dataset. Column 2 uses the Tokyo City Statistics Division (1926a, pp. 16; 244–245); Tokyo City Office (1921, pp. 726–757).
@@
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) focusing on factories with more than or equal to 15 workers show that roughly 90% of male workers had been engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.363636 I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in column 2.
@@
The official report of the Tsukishima Survey suggests that roughly nine out of ten factory workers in the skilled-age range were classified as skilled workers. Roughly six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^38] Online Appendix A.5 summarizes this point in detail using the complete survey on the primary school students in 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^39] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector will constitute the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is 33, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^40] Specific occupations are not available for most heads because they usually list “factory workers” in the occupation section. However, sub-categories available at a few heads indicate the fact that they were lathe operator and finisher, the typical skilled workers in machinery factories. Online Appendix A.6 summarizes the works in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $53$ yen. In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($55$ yen). [^41] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30-40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima are materially similar. To summarize, this evidence indicates that the THBS sample is useful in inferring the mean tendency of those households.

### v7

@@ Abstract
I analyze the risk-coping behaviors among factory worker households in early 20th-century Tokyo. I have digitized a unique daily longitudinal survey dataset on household budgets to analyze how consumption is impacted by idiosyncratic shocks. I find that while the households were so vulnerable that the shocks impacted their consumption levels, the income elasticity for food consumption is relatively low in the short-run perspective. The result from mechanism analysis suggests that credit purchases with local retailers smoothed short-run food consumption. The event-study analysis using the adverse health shock as the idiosyncratic income shock confirms the robustness of the results. I also find evidence that the misassignment of payday in data aggregation results in a systematic attenuation bias due to measurement error in the standard consumption smoothing regressions.
@@ 2.1 Japanese Economy around the First World War
Note: This figure shows the gross national product per capita (solid line) and personal consumption expenditure (dashed line) between 1910 and 1926. The prices are deflated using the CPI in 1934-1935 price.
Sources: Created by the author using Ohkawa et. al. (1974, p. 237).
@@ 2.2 Tsukishima: A Representative Manufacturing Area
Notes: The name in each lattice indicates the administrative ward in Tokyo. The gray ward indicates the Kyobashi Ward and an island in the Kyobashi Ward indicates Tsukishima. The border surrounded by Asakusa, Honjyo, Nihonbashi, and Fukagawa wards includes the Sumida River. The borders of each ward are based on the administrative district in 1920. Geographical coordinate system is based on JGD2000/(B, L).
Source: Created by the author using the official shapefile (Ministry of Land, Infrastructure, Transport and Tourism, database).
@@
Notes: Blocks in green and light green indicate the shipyards. Blocks in red and pink indicate the factories in the machinery sector. Blocks in blue and light blue indicate the factories in the other manufacturing sectors. Blocks in brown are the warehouses. Small circles colored in black and white show the workplaces in the machinery sector and the other manufacturing sectors, respectively.
Source: Department of Health, Ministry of the Interior 1923c, second map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@
Notes:
Panel A: This panel summarizes the industrial structures based on the occupations of male workers measured in the 1920 population census. Each share is calculated as the number of males worked as regular workers in each sector divided by the total number of male workers (in percentage). All the figures do not include the number of unemployed (mugyō sha) males.
Panel B: The average household size is the number of people divided by the number of households. The share of married males is the number of married males living with/without spouses divided by the total number of males (in percentage). The average age of males is calculated using the tables for population by age group reported in the census. Since the tables are open-ended, the class for over 60 years old is rounded as $64.5$ following the range of the former category (i.e., 50–59). All the figures in this panel do not include a small number of quasi-households (jyun setai), which include the person, such as the students in dormitories and the patients in hospitals.
The figures for Tsukishima listed in column 3 of both panels are based on the total sum of statistics of all the blocks (chōme) in Tsukishima (i.e., from Tsukudajima to Tsukishima dōri jyūni chōme).
Sources: Columns 1 and 2 of Panel A use the Statistics Bureau of the Cabinet (1929c, pp. 18–19). Columns 1 and 2 of Panel B use the Statistics Bureau of the Cabinet (1929c, pp. 38–43); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 3 of Panel A uses the Tokyo City Office (1922c, pp. 42–47). Column 3 of Panel B uses the Tokyo City Office (1922a, pp. 26–31); Tokyo City Office (1922b, pp. 2–3; 42–46).
@@ 2.3 Risk-Coping Institutions
Working-class households had been exposed to several risks that led to losses of income at that time. Generally, social insurance, such as health, accident, and unemployment insurance, can help mitigate the idiosyncratic income shocks on the workers. [^24] Gertler and Gruber (2002) shows the impacts of illness on consumption smoothing behavior in early 1990s Indonesia and argues the importance of public support for the illness. Kantor and Fishback (1996) show that the installation of workers’ compensation reduced the precautionary savings in the interwar U.S.. Gruber (1997) provides evidence that unemployment insurance smooths individual consumption as the case of the U.S. from 1968–1987. However, comprehensive social insurance systems did not exist in the early 1920s in Japan. The former health insurance was enacted in 1927, and unemployment insurance had not been established throughout the prewar period. In addition, the Factory Act enacted in 1916 did also not specify compensation for the sickness of workers. Therefore, the workers had basically no public support for their illness, injuries, and joblessness. Instead, they might have used private financial institutions to deal with the idiosyncratic risks.
@@
Several financial institutions were indeed available for working-class households. A representative savings institution for households was the postal savings (yūbin chokin). In Tokyo City, roughly $1.4$ million people had postal savings accounts, covering approximately $67$% of the entire citizens in 1920 (Online Appendix Table A.2). The average savings amount per capita in Kyobashi Ward was $44.5$ yen (Tokyo City Office 1921, pp. 892–893), which is less than one month’s earnings of the skilled-factory workers (Section 3.2). Savings bank (chochiku ginkō) is another famous savings institution among the working class in the 1910s (Ito and Saito 2019, pp. 79–81). The number of people with savings accounts was roughly $1.4$ million, accounting for approximately $58$% of total people in Tokyo (Online Appendix Table A.2). However, $84.8$% of all the depositors were the commerce and miscellaneous workers, and the manufacturing sector shared only $11.5$% in Kyobashi Ward (Online Appendix Table A.3). This reflects the fact that savings bank was an institution designed for lower-income workers than factory workers (Tokyo Institute for Municipal Research 1925b, p. 81). This is consistent with the fact that Tsukishima did not have any savings bank branches, while there were two postal offices (Section 2.2). Although another savings institution was the mutual loan association (mujin), it had been used among small business owners and merchants for running their businesses (Online Appendix A.2). [^25] Generally, the ordinary bank in prewar Japan was the institution for providing large loans to businesses (Teranishi 2011). There was no bank in Tsukishima, as workers did not borrow money from banks to make ends meet. The official report states that “a commercial bank is not necessary in a laboring area such as this Tsukishima, and even for large entrepreneurs such as factory owners, there is no particular need to establish a bank for their business transactions, and transactions with banks in other areas are sufficient” (Department of Health, Ministry of the Interior 1923a, p. 48).
@@
There also be a few lending institutions. Credit purchases (kakegai) in retail shops were the most popular institution. Although the systematic statistics on credit purchases are unavailable, the THBS households often used credit purchases in various retail shops for rice, fish, vegetables, firewood, and charcoal. [^26] Cooperative societies were also available for worker households to buy daily commodities, albeit they were not lending institutions. There were 26 cooperative societies in Tokyo prefecture in 1924, of which 23 were organized by the citizens and workers (Central Federation of Industrial Associations 1925, p. 57–59). However, while the two largest societies had roughly $4,000$ members, the others were generally small societies with less than 100 members. This figure accounts for only a few percentage points in the workers in the city, given that the number of male workers was more than 768 thousand. Most of the THBS households had not used the cooperatives. I will describe this institution in detail in Section 5. The pawnshop (shichiya) was another lending institution. The accessibilities were higher than the other lending institutions: [^27] The number of pawnshops in Tokyo were $1,334$ in 1918 and $1,261$ in 1919, respectively. The total number of cases was $8,226,883$ and $7,573,406$ each year. These figures are more than three times greater than the entire population in Tokyo City. there were $13$ pawnshops in Tsukishima, more than twice the large number of moneylenders (Section 2.2). However, the core users of the pawnshops were the low-income working-class households rather than the factory workers (Tokyo City Social Affairs Bureau 1921, p. 9). According to the statistics on the pawnshop users in Tokyo prefecture in 1923, the most representative users were the day laborers and workers classified as “miscellaneous”, accounting for $63.4$% of all users. In contrast, those for agriculture, commerce, and manufacturing sectors were only $1.7$, $19.5$, and $15.6$%, respectively (Tokyo Institute for Municipal Research 1926, p. 25). The inexpensive article means that pawnshops were used to deal with very short-run necessities only. Among the pawnshops of Kyobashi, for example, the average amount per case was $7.8$ yen, which accounts for $12$% of the monthly disposable income of the THBS households (Panel A of Table 3). In this light, the pawnshop is a small lending institution for low-income working-class households who might have faced credit constraints. Lenders did not need to screen the borrower’s credit, and the borrowers did not worry about incurring heavy debt as the main articles were inexpensive clothes (Shibuya et al. 1982). In addition, the interest rates were regulated by the Pawnbroker Regulation Act of 1895, and the average redemption rate was substantially high under the lower interest rates. [^28] This means that the average redemption rate was substantially high: it was approximately $94$% in 1920 among $88$ pawnshops in Kyobashi ($401,222$ cases/$427,265$ cases) (Tokyo City Office 1922b, pp. 888–889).
@@
Finally, life insurance was also not designed to compensate for temporary income losses. Generally, life insurance companies provided death insurance, meaning that temporary income reductions were not covered by insurance. Moreover, they did not target the working-class as customers (Usami 1984, pp. 87–88). Although postal life insurance (kani seimei hoken) operated by the government started in 1916, it only provided whole life insurance and endowment insurance. Thus, it also never covered the health shocks nor layoffs (Usami 1984, pp. 102–107).  [^29] There were also several employment agencies that introduced jobs to low-income job-seekers in Tokyo. However, a large part of them was unscrupulous and not commonly used at that time (Machida 2016).
To summarize, while there were several institutions, the available institutions for the factory workers were few in the practical sense. For savings institutions, the postal savings might have been the best available device for Tsukishima workers. Regarding lending institutions, credit purchase is the plausible device among factory workers. While pawnshops were accessible financial institutions, they might not provide sufficient money as the articles were usually inexpensive. The factory workers in Tsukishima might have combined these devices to deal with the idiosyncratic income shocks.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption smoothing and risk-coping behaviors requires the panel structures for aggregate income and expenditure (Section 4). Among the 40 THBS households, there are five cross-sectional units in our semi-monthly data. [^35] Section 3.3 provides details on the definitions of aggregate income and expenditure. I excluded these five cross-sectional units, leaving 35 THBS households. Next, a household with incomplete income information, which is necessary for my analysis, is removed. Finally, I exclude a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series includes 33 households with panel structures from 12th January to 11th November 1919.363636I use adjusted monthly panel dataset for subanalysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset includes 26 households for the same period. Note that several units have only three semi-month cells, say less than two adjusted months. These units have panel structure in the semi-monthly panel, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
Before assessing the representatives of the THBS sample, I test the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null results. This result is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference is less likely to be correlated with unbalancedness.
@@
Average monthly earnings 59.1 (median = 56.2) 56.0
95%CI [53.7, 64.5]
@@
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. Columns 2 and 3 use the classification definition in the Statistics of the National Police Agency (column 1). Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be regarded as one of the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389).
Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The third to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5).
Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Of the 33 THBS households, information from 26 households for which adjusted monthly income can be traced back is used in this panel to follow the unit in the manufacturing census. Comparison based on the average semi-monthly income from all 33 households shows a similar result. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30-40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation.
Sources: Panel A: Column 1 uses the National Police Agency (1920b, p. 236). Column 2 uses the Department of Health, Ministry of the Interior (1923a, p. 385–386). Column 3 uses the Department of Health, Ministry of the Interior (1923a, p. 154). Panel B: Column 1 uses the Tokyo City Office (1922a, pp. 26–31; 262–283); Tokyo City Office (1922b, pp. 2–3; 42–46). Column 2 uses the National Police Agency (1920b, p. 53). The figures in column 3 are calculated by the author using the THBS dataset. Panel C: Column 1 uses the THBS dataset. Column 2 uses the Tokyo City Statistics Division (1926a, pp. 16; 244–245); Tokyo City Office (1921, pp. 726–757).
@@
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) focusing on factories with more than or equal to 15 workers show that roughly 90% of male workers had been engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.373737 I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in column 2.
@@
The official report of the Tsukishima Survey suggests that roughly nine out of ten factory workers in the skilled-age range were classified as skilled workers. Roughly six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^39] Online Appendix A.5 summarizes this point in detail using the complete survey on the primary school students in 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^40] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector will constitute the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is 33, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^41] Specific occupations are not available for most heads because they usually list “factory workers” in the occupation section. However, sub-categories available at a few heads indicate the fact that they were lathe operator and finisher, the typical skilled workers in machinery factories. Online Appendix A.6 summarizes the works in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen). In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen). [^42] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30-40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima are materially similar. To summarize, this evidence indicates that the THBS sample is useful in inferring the mean tendency of those households.